### v2

# Improving User Mental Models of XAI Systems with an Inclusive Design Approach

Abstract.
Explainable Artificial Intelligence (XAI) systems aim to improve users’ understanding of AI but rarely consider the inclusivity aspects of XAI. Without inclusive approaches, improving explanations might not work well for everyone. This study investigates leveraging users’ diverse problem-solving styles as an inclusive strategy to fix an XAI prototype, with the ultimate goal of improving users’ mental models of AI. We ran a between-subject study with 69 participants. Our results show that the inclusivity fixes increased participants’ engagement with explanations and produced significantly improved mental models. Analyzing differences in mental model scores further highlighted specific inclusivity fixes that contributed to the significant improvement in the mental model. To our surprise, the inclusivity fixes did not improve participants’ prediction performance. However, the fixes did improve inclusivity support for women and promoted equity by reducing the gender gap.
Diversity, Inclusive AI, Explainable AI, Mental Model
copyright: acmcopyrightjournalyear: $9$doi: 99.9999/9999999.9999999journal: TIISpublicationmonth: 99ccs: Human-centered computing Empirical studies in HCI
1. Introduction
“The curb-cut effect underscores the foundational belief that we are one nation, that we rise or fall together. Without equity, there can be neither progress nor prosperity.” (Blackwell, 2016) —Angela Glover Blackwell, author of the book “Curb-Cut Effect”.
Angela Glover Blackwell’s thoughts on equity resonate with our focus on inclusive XAI. While Blackwell talks about equity, we’re focusing on inclusivity in this paper. As researchers are working to make XAI better, inclusivity asks the question “better for whom?”
By “better”, we mean promoting better mental models of how a user’s AI system is reasoning. In XAI, mental model is defined as the user’s understanding of how an AI system works (Anderson et al., 2020). More formally, mental models are “internal representations that people build based on their experiences in the real world” (Norman and Gentner, 1983). Helping users form suitable mental models is the aim of XAI systems. Although such XAI systems are gaining attention as AI becomes more widespread, the challenges of understanding these systems are also on the rise (Weld and Bansal, 2019).
Improving explanations for one specific user group may not always benefit others if we don’t adopt inclusive approaches in existing AI development practices. One particular inclusive approach involves recognizing users’ diverse problem-solving styles. While demographic diversity is important, considering problem-solving diversity generates actionable design directions (Hamid et al., 2024). Several studies showed how addressing users’ diverse problem-solving approaches reduced inclusivity biases from non-AI applications and systems (Vorvoreanu et al., 2019; Guizani et al., 2022; Burnett et al., 2016). This paper investigates whether focusing on users’ diverse problem-solving styles can reduce inclusivity issues across an information-rich XAI system, ultimately aiming to improve users’ mental models.
To do so, we evaluated how well a group of end users fared in two versions of an XAI prototype. These versions were the work of a team of AI practitioners to address inclusivity issues within their XAI prototype. Their prototype showed a modified version of Tic-Tac-Toe games played sequentially by two AI agents. The AI practitioners leveraged the GenderMag inclusivity method (Hamid et al., 2024; Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019) to find and fix the inclusivity issues. We ran a between-subject study where a group of 34 participants used the “Original” version and 35 participants used the “Post-GenderMag” version. We studied participants’ predictions, their ability to detect flaws in the AI agents, their perception of the AI agents’ behavior, and lastly their mental models. Our investigation centered around these three research questions:
  • RQ1: Did mental models improve with the PostGenderMag version?
  • RQ2: Did prediction performance improve with the PostGenderMag version?
  • RQ3: Did the Post-GenderMag version provide better inclusivity support?
2. Related Works and Background
In recent years, Human-AI Interaction (HAI) research has focused on user performance, system reliability, trustworthiness, and safety (Shneiderman, 2020), while also exploring the dynamics of human and AI control levels (Shneiderman, 2022). The growing body of HAI works aims to increase transparency (Rader et al., 2018), explain AI behaviors (Herlocker et al., 2000; Kulesza et al., 2015), provide guidelines for AI practitioners (Amershi et al., 2019), and address ethical principles in governance (Shneiderman, 2021). Additionally, HAI research advocates for human-centric concerns such as well-being, privacy, and cognitive capacities in the adoption of AI technologies (Ozmen Garibay et al., 2023; Shneiderman, 2022). In our study, we are specifically interested in bringing diverse users’ problem-solving styles together with HAI to improve XAI’s inclusivity.
2.1. Problem-Solving Style and HAI/XAI
There’s been limited investigation into how users’ problem-solving styles influence HAI/XAI. Although AI researchers have examined users’ explanation preferences, they’ve largely overlooked individual differences within these preferences (Dodge et al., 2019). Some research used users’ cognitive diversity to understand why users behave or judge differently when interacting with AI systems. For example, when looking at how fair end-users’ perception of AI was, Dodge et al. (Dodge et al., 2019) assessed the impacts of two factors: i) four types of explanations and ii) data processing. Their findings revealed that participants’ need for cognition significantly impacted their perception of the AI’s fairness. Similarly, Millecamp et al.’s work (Millecamp et al., 2020) found that diverse end-users’ cognitive styles showed differences in their trends to interact with the AI’s explanations. This investigation discussed that the variation in the definitions of cognitive styles did not have much effect and any of the definition variations could help identify valuable insights.
Goldberg et al. (Goldberg, 1990) presented an alternative approach to capturing individual differences through five personality traits— Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (OCEAN) (Goldberg, 1990). They found that one of the personality traits, “Openness” influenced users’ interaction with the interface. Jiang et al. (Jiang et al., 2000) also used a personality trait, self-confidence to evaluate participants’ acceptance of the expert system’s output. Other researchers also used some form of problem-solving styles, cognitive styles, and personality traits to explain user behavior (Giboney et al., 2012; Kulesza et al., 2012; Felmingham et al., 2021; Djulbegovic et al., 2014). However, none of these works considered a wider range of problem-solving styles as we did in our study.
In this study, we used problem-solving styles defined by users’— attitude towards risk, computer self-efficacy, motivations, information processing style, and learning style.
2.2. GenderMag
We used the GenderMag method to define the end users’ five problem-solving style ranges used in this study. GenderMag, also known as the Gender-Inclusiveness Magnifier, is a usability inspection method to find and fix gender-inclusivity bugs in problem-solving software (Hamid et al., 2024; Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019). This method focuses on five facets of gender differences (statistical), brings them to life through three faceted personas, and encapsulates the use of these facets into a systematic process through a gender specialization of the Cognitive Walkthrough (CW) (Wharton et al., 1994). Software practitioners have used this method to evaluate the software they design and develop from a gender-inclusiveness perspective (Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019).
The GenderMag method’s five facets capture five factors in an individual’s problem-solving style: their motivations for using technology, information processing styles, computer self-efficacy, attitude toward risk in using unfamiliar technology features, and learning style (by process vs. by tinkering) (Burnett et al., 2016; Hamid et al., 2024). The “inclusivity bugs” uncovered by GenderMag are instances where a technology product fails to fully accommodate the complete range of values associated with the five facets, leading to a disproportionate impact on individuals whose problem-solving styles are unsupported. These bugs are also considered gender-inclusivity issues because the facets reflect statistical differences among gender preferences in problem-solving styles (Burnett et al., 2016; Guizani et al., 2022; Vorvoreanu et al., 2019; Hamid et al., 2024). GenderMag reveals inclusivity issues that are statistically associated with gender preferences. GenderMag does not use peoples’ gender identities to find inclusivity concerns, it uses only their problem-solving styles.
The GenderMag method has three personas: Abi (Abigail/Abishek), Pat (Patricia/Patrick), and Tim (Timara/Timothy) (see Table 1). Statistically, the problem-solving styles of Abi disproportionately skew toward women, for Tim they disproportionately skew toward men, and Pat provides a third set of values within Abi and Tim (Burnett et al., 2016; Hamid et al., 2024).
Table 1. Three GenderMag personas and their problem-solving styles
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Abigail/Abishek (”Abi”) Patricia/Patrick (”Pat”) Timara/Timothy (”Tim”)
Attitude Toward Risk Risk-Averse Risk-Averse Risk-Tolerant
Self-Efficacy Lower Medium Higher
Motivations Task-Oriented Task-Oriented Tech-Oriented
Information Processing Style Comprehensive Comprehensive Selective
Learning: by Process vs. by Tinkering Process-Oriented Learns by tinkering: tinkers reflectively Learns by tinkering (sometimes to excess)
Abi and Tim have problem-solving styles at each facet’s endpoints, and Pat has a mix of possible facet values.
The GenderMag faceted personas are integrated into a systematic specialized Cognitive Walkthrough process. A group of evaluators adopt a specific GenderMag persona, say Abi, customize the persona’s background (optional), and systematically “walkthrough” each step of a use case. At each step, they answer three questions: i) will ¡Abi¿ have formed this subgoal as a step to their overall goal? why?, ii) will ¡Abi¿ take the action the developers expected at this step? Why?, and iii) if ¡Abi¿ took the action, will they know that they did the right thing and are making progress toward their goal? Why? If the evaluators find any answer is negative because of Abi’s particular problem-solving styles, they report an inclusivity bug, provide their reasons, and note down the problem-solving styles that support their reasonings. Later, when fixing an inclusivity bug, these notes ensure that the fixes are designed around Abi’s problem-solving styles that highlighted the bug.
The evaluators relevant to our study were a group of AI practitioners whom we called “Team Game” in this paper. Team Game evaluated their own XAI prototype from the perspective of the Abi persona. They chose the Abi persona as Abi represents the group of users who are most often left out of the development process which results in more barriers for these users. Team Game’s before GenderMag and after GenderMag version of the prototype were then used in our study.
2.3. XAI in Sequential Domains
The AI practitioners of our study designed their XAI prototype specifically for a sequential domain. A sequential domain refers to a specific situation or area where actions or events unfold in a particular order, following a series of steps or stages in a predetermined sequence. For instance, consider a board game like Chess, where players take turns making moves and the progression of the game follows an ordered sequence of moves. In recent years, there has been a growing interest in developing AI systems that can play sequential games like Chess, Shogi, and Go. This has led to the development of AI players that can compete with and even outperform the world champions of these games (Silver et al., 2016, 2018). Real-Time Strategy (RTS) games have also seen similar achievements (Vinyals et al., 2019b, a).
Despite such progress, explaining AI players’ actions in sequential domains remains a major challenge. In an attempt to understand this challenge, Metoyer et al. (Metoyer et al., 2010) studied how experienced users explain strategies and gameplay in RTS games. In another study, Penney et al. (Penney et al., 2018) found that even experienced users encounter challenges when seeking information about an intelligent agent in an RTS setting. One result was that participants showed a preference for “What” information over “Why”. In a sequential game domain, understanding the causal relationships among relevant variables within the game can also help users (Madumal et al., 2020). As Madumal et al. (Madumal et al., 2020) showed offering causal explanations of AI agents can enhance participants’ ability to predict tasks, as well as increase their satisfaction and trust. In another study, Dodge et al. (Dodge et al., 2022) observed that when ranking AI agents playing a sequential game like Tic-Tac-Toe, participants highlighted the importance of detailed explanations to differentiate between closely matched agents. That study also noted that participants showed a preference for receiving multiple explanations over a single one when making decisions.
2.4. Mental Models and Explainable AI (XAI)
The opaque nature of AI systems poses a challenge for end-users with low or no expertise in AI (Wortham et al., 2017; Tsai and Carroll, 2022). This especially hinders users’ ability to seamlessly interpret and trust the system’s output (Kulesza et al., 2015). Lim et. al (Lim et al., 2009) regarded such lack of clarity as an intelligibility concern for AI systems.
To improve the clarity of an AI system, researchers and developers often provide explanations of the system to its end-users (Alipour et al., 2021; Brachman et al., 2023; Schraagen et al., 2020; Pierson et al., 2024). According to Hoffman et al., explanations that are effective and satisfactory for users can facilitate the development of a good mental model (Hoffman et al., 2019). Kulesza et al. demonstrated that providing explanations about the debugging process of a machine-learning system improved participants’ understanding of the system (Kulesza et al., 2015, 2010). Also, the explanations improved their participants’ ability to rectify errors. Other studies also found additional benefits of adding explanations such as improvements in users’ self-reported satisfaction (Kapoor et al., 2010; Amershi et al., 2014) and feelings of trust towards the AI systems (Lim et al., 2009; Tsai et al., 2021; Schraagen et al., 2020).
Mental models have been captured using different methods including questionnaires about the agent, mental model drawings, and task performance (Brachman et al., 2023; Schraagen et al., 2020; Zhang et al., 2024; Bansal et al., 2019; Alipour et al., 2021). Hoffman et al. (Hoffman et al., 2019) proposed measures that evaluate the quality, user satisfaction, and user understanding of the explanations. According to Hoffman et al., mental models can be deduced from empirical evidence, even though individuals might find it challenging to fully articulate their understanding (Hoffman et al., 2019). However, the authors suggested that through guided task reflection, people can effectively communicate their understanding of an event or system, including its fundamental concepts and principles. The authors provided a list of 11 methods that researchers have used to derive mental models across various domains, including cognitive psychology, expert system knowledge acquisition, instructional design, and educational computing. From these 11 methods, Anderson et al. used the “Prediction Task” and “Task Reflection” to understand users’ mental models of a Reinforcement Learning agent (Anderson et al., 2019). In our paper, we have utilized these two methods and also integrated the “Glitch Detector Task”, and “Self-Explanation Task” from Hoffman et al.’s 11 methods to know if users were able to find any flaws in the given explanations and their own understanding of the AI agent.
Thus, although HAI and XAI research have made significant progress in exploring users’ mental models, they often overlooked the potential impact of users’ diverse problem-solving styles on their overall understanding of AI systems. Recognizing this gap, we wanted to check the effectiveness of GenderMag-driven cognitive walkthroughs that put users’ diverse problem-solving styles at the center stage of finding and fixing inclusivity issues. Historically, HCI practitioners leveraged GenderMag for a wide variety of user experience improvements such as inclusivity bug removal (Guizani et al., 2022), task performance improvement (Vorvoreanu et al., 2019), customer rating improvement (Burnett et al., 2017), etc. However, no prior GenderMag research has evaluated GenderMag’s applicability and effectiveness within AI-powered systems. Addressing this gap in GenderMag studies, our goal is to investigate how successful GenderMag fixes can be in improving users’ engagement and their mental models of AI.
3. Methodology
3.1. Experimental Design
We set up a between-subject study in a lab to evaluate the two versions of Team Game’s XAI prototype.
3.1.1. Participants
To recruit participants, we distributed flyers through our university’s internal mailing lists and posted them around campus. The recruitment emails and flyers had the invitation link to a Qualtrics survey, where we collected demographic information and asked two eligibility questions to make sure our participants were at least 18 years of age and had not taken any AI courses as we wanted to recruit people with no AI background. Based on these criteria, we dropped 41 responses and sent out emails to the remaining 110 eligible respondents individually. Ultimately, 77 respondents accepted the invitations and 69 participated in our study. These 69 participants came from diverse backgrounds, spanning different genders, age groups, and college statuses (all summarized in Table 2), as well as 39 unique college majors.
Table 2. Summary of the 69 participants’ demographics
Gender #
Woman 36
Man 28
Non-Binary 2
Prefer not to state 2
Woman, Transgender 1
Age #
18-19 9
20-29 38
30-39 17
40-49 2
50-59 3
College Status #
Undergraduate 31
Graduate 21
Has graduated 9
Other 8
3.1.2. Procedures
We first assigned participants to sessions based on their availability and then randomly assigned each session to one of two versions of Team Game’s game prototype: the “Original” and the “Post-GenderMag”. The Original sessions included 34 participants, while the Post-GenderMag sessions had 35 participants. Each session lasted 2 hours and had 2 to 6 participants. We conducted a total of 17 sessions from August 18 to August 28, 2023.
Upon arriving at the lab, participants filled out the GenderMag survey (provided in the supplemental document of (Vorvoreanu et al., 2019)) in which participants answered Likert-scale items about their problem-solving styles. Then we began the session with a 25-minute hands-on tutorial covering the game, agents, explanations, and participants’ tasks. Afterward, participants went through the first three moves of a practice game to get familiar with the prototype and the study questions.
In the main part of the study session, participants observed 3 games of Tic-Tac-Toe played by 2 AI agents: Agent Blue (X) and Agent Red (O). Before Agent Blue made a move, participants predicted the move, provided reasoning, and their confidence in it. After the move, they identified flaws and suggested changes. After each game, they answered two After-Action Review for AI (AAR/AI) questions. Once all three games concluded, they completed post-task questions covering agent descriptions, explanation feedback, and subjective workload assessment using NASA TLX (NASA, [n. d.]). Table 3 lists all the questions.
Table 3. The questions we asked both groups of participants
Before Agent Blue made a move
- Which square do you think the blue agent will place its next move?
- Why do you think that the agent will place its move there? (A few sentences, as detailed as possible)
- How confident are you that the agent will select the square that you predicted? On a scale of 1 (not confident at all) to 10 (extremely confident).
After Agent Blue made a move
- From looking at the explanations (right), do you think there is a flaw in the agent’s reasoning? If so, what is that flaw? (A few sentences, as detailed as possible)
- What CHANGES would you make in the decisions taken by the agent? (A few sentences, as detailed as possible.)
After each game: AAR/AI
- What happened in this game (write any good, bad, or interesting things you’ve observed in these past moves and/or games)? (A few sentences, as detailed as possible.)
- Is there anything in the explanation that helps you understand why the AI you’re assessing did the things it did? Please specify which explanation you are referring to. ($\sim$2 sentences)
After finishing all 3 games
- Please describe the steps in the agent’s approach or method. (A few sentences, as detailed as possible.)
- Explanation feedback  [^1] Participants provided feedback for each of the three explanations, BTW, STT, and OTB. They responded on a 7-point Likert-scale from Very Strongly Disagree to Strongly Agree.: i) the explanation helped in understanding Agent Blue, and ii) the explanation was sufficiently detailed.
- NASA TLX.
From these questions, we have not used the responses from “Please describe the steps in the agent’s approach or method…”. When the user study was over, we found that Agent Blue’s noise levels were different between each game. This means participants would be describing multiple agents rather than one. Despite Team Game’s lead (who created the agent) suggesting that the noise differences didn’t affect the agent’s behavior much, we excluded these responses from our analysis to ensure clarity in our evaluation of users’ understanding of AI.
3.1.3. Qualitative Analysis Methods
After all the study sessions were over, we used affinity diagramming to organize the participant responses into concepts participants expressed. Five of the researchers independently marked participants’ responses that described some aspect of their mental model of the AI agents. We then discussed the results as a group and developed the initial concept codeset. We also met with Team Game’s lead (the subject matter expert who had created the AI agents) to further modify the initial concept codeset. Finally, the concept codeset had 35 concept codes divided into four groups: “Prediction Why”, “Strategy Should Be”, “Flaws”, and “HowAgentBehaves” (Table LABEL:tab:codes_definitions). Four of the researchers (R1, R2, R3, and R4) then worked in pairs of two as per Table 4 to independently code ¿=20% of each of the four concept code groups. Table 4 lists the inter-rater reliability rates (IRR) (Stemler, 2004) of the agreement tests. Given the high IRRs, one researcher from each coder pair finished the rest of the coding for their assigned concept code group.
Table 4. Interrater-reliability results (IRR) for each concept code group
Concept Code Group IRRTest DataCoders
Prediction Why 82% 20% R1, R2
Flaws 80% 26% R1, R3
Strategy Should be 81% 25% R1, R2
HowAgentBehaves 81% 20% R1, R4
3.2. Where the Two Versions of the Prototype Came From
Before our study, Team Game used GenderMag to find and fix inclusivity issues in their game prototype. During the “find” sessions, Team Game identified inclusivity issues within the Original prototype (Figure 1). In the “fix” sessions, they proposed design solutions to potentially address these issues. These sessions took place over a period of five weeks via Zoom.
Figure 1. The Original prototype during a game. Columns are labeled A-I and rows are labeled 1-4. Squares can be identified using their column name followed by the row number (e.g., E1). The game progression is controlled by clicking the prior move and next move buttons below the gameboard.
Team Game started the find sessions by customizing the Abi persona, which they then used during the GenderMag walkthrough to identify inclusivity issues within their prototype. They identified issues that spanned all of GenderMag’s five problem-solving styles outlined in Table 1. Team Game engaged in find sessions totaling about five hours.
In the fix sessions, Team Game systematically reviewed all identified inclusivity issues from the find sessions spending about eight hours in total. Team Game devised 15 proposed fixes during these sessions. Table 5 presents two of these fixes which we later utilized to illustrate our study findings. All the 15 fixes are listed in Table LABEL:tab:fixes_all in the appendix.
Table 5. Two fixes, why Team Game fixed them (problem-solving styles impacted), and the changes made for the fixes
Fix-ID & Fix Name Fixed-Why + Problem Solving Styles(PSS) Changes Made for the Fixes
Fix-2: Exact values and Exact actions “She is not clear on what she should do…..will not take this step.” PSS: Risk, Learn, Info, SE, Motiv. – Showed exact win%, loss%, and drawfor every rectangle in the BTW, STT, and OTB explanations using tooltips. – Added tooltips that informed about the exact actions offered by the next and the prior move buttons.
Fix-7: Top 5 Moves “…but they may not be sure how to reason with this information.” PSS: Risk, Learn, Info, Motiv. – Added “Top 5” moves with their score, win, loss, and draw percentages.
Rest of the Fixes are in Table LABEL:tab:fixes_all.
We then implemented 13 of Team Game’s 15 fixes. We could not implement Fix-14 and Fix-15 due to their technical difficulties and our time constraints. Figure 2 shows the Post-GenderMag version of Team Game’s prototype.
Figure 2. The Post-GenderMag prototype during a game. Some fixes resulted in new additions such as the (A) Game History, (B) Game Log, and (C) Top 5 Moves. These are a few of the features that differentiate the Post-GenderMag version at this state of the game from the Original version (Figure 1) at the same state.
We call the original version “Original” and the redesigned version “Post-GenderMag” as its changes came from the GenderMag-driven cognitive walkthroughs. With these two versions, we aimed to address our research questions about Team Game’s GenderMag-driven fixes’ effects on participants’ mental models, prediction performance, and explanation ratings.
3.3. Game Overview and Prototype
Team Game was working in the sequential decision domain of MNK games, based on Dodge et al.’s original prototype (Dodge et al., 2022). MNK games are an abstraction of Tic-Tac-Toe where M and N represent the dimensions of the gameboard ($M\times N$) and the goal is to get a sequence of length K. To simplify the prototype presented in Dodge et al. (Dodge et al., 2022) and enable Team Game to make changes, we re-created the original prototype in a web-based format using Svelte (JavaScript framework) seen in Figure 1. In this MNK game, the gameboard—on the left half of the prototype—has 9 columns and 4 rows (36 squares) where the winning condition is getting 4 in-a-row (9-4-4). The game is played by two AI agents, Agent Blue (X) and Agent Red (O), that use a convolutional neural network to predict outcome tuples $O=(Win\%,Loss\%,Draw\%)$ for each square on the gameboard before making a move. Each square is then assigned a “Score” using the function $Win\%-Loss\%$. The agents make moves based on the scores. To the right of the gameboard, the prototype provides three explanations for Agent Blue’s move but not for Agent Red. This prototype had the same explanations presented in Dodge et al. (Dodge et al., 2022), namely Scores Best-To-Worst (BTW), Scores Through Time (STT), and Scores On-the-Board (OTB).
3.3.1. Explanation Type: Scores Best-to-Worst (BTW)
The Scores Best-to-Worst (BTW) explanation presents all of Agent Blue’s predicted scores for the 36 gameboard squares ranked from the best move to the worst move on a graph (Figure 3). The X-axis is the order of scores, from the best move on the left to the worst moves on the right and the Y-axis represents the score from -100% to 100%. Agent Blue assigns very low scores for the illegal moves (moves that are already taken). After each move by Agent Blue, a new series of rectangles will appear in the explanation and the prior set of scores will get a lighter color. The most recent data series will have the darkest color and can be interacted with by hovering. Hovering over a score in the most recent data series will highlight its corresponding square on the gameboard and its past scores in all the explanations. Hovering over a square on the gameboard will highlight its corresponding scores in BTW (and also in other 2 explanations) over the past moves.
Figure 3. BTW in the Original prototype. Three data series of scores by Agent Blue after the agent made its third move. The blue data series is the most recent scores the agent calculated.
3.3.2. Explanation Type: Scores Through Time (STT)
The Scores Through Time (STT) explanation shows how the distribution of scores changes throughout the game whenever Agent Blue makes a move (Figure 4). The X-axis is time (decision points) and the Y-axis is the score. When Agent Blue makes a move, a new column of scores appears. The score with the maximum score in the column is highlighted in a darker color for the current move. Hovering over a score in the column highlights its corresponding square in the gameboard and the score of that square in prior decisions. Like the prior explanation BTW, only the most recent series of scores can be hovered over. Hovering over a gameboard square will highlight its score for every decision through time.
Figure 4. STT in the Original prototype. Agent Blue has played 3 moves so there are 3 columns and the number under each column represents the move number.
3.3.3. Explanation Type: Scores On-the-Board (OTB)
The last explanation, Scores On-the-Board (OTB), shows the scores through time at the level of each square where each square has its own version of STT (Figure 5). Each little dot shows Agent Blue’s evaluation of that square over different moves or decision points. Squares that have been used for moves have a highlighted border in the same color (blue or red) as the agent that picked the move. Hovering over a square in the gameboard highlights the corresponding chart in the explanation and vice versa. A new data point is added to each square’s chart every time Agent Blue makes a move.
Figure 5. OTB in the Original prototype. The square highlighted in yellow is the most recent move by Agent Blue which has a high score.
3.4. Measuring Mental Models for this Game
3.4.1. Mental Model Rubric
We needed a mental model rubric to calculate the mental model score of each participant. From the four concept code groups of the initial codeset, Team Game’s lead, who created the AI agents, selected two groups, “Flaws” and “HowAgentBehaves” for the mental model rubric (Table 6).
“Flaws”: The agent creator picked the following three concepts under “Flaws” for the rubric:
– Square X is miscalculated.
– Agent miss selection.
– Agent chose the wrong action.
The agent creator pointed out that to identify flaws, participants needed to understand the scoring system, the agent’s decision-making process, and the gameboard. They also mentioned both local and global understanding of the Agents’ moves were necessary for detecting flaws. For these reasons, the agent creator picked the above three concepts that reflected Agent Blue’s decision-making “Flaws”. They assigned positive weights to these concepts in the mental model rubric.
“HowAgentBehaves”: The agent creator picked seven concepts from the “HowAgentBehaves” group for the mental model rubric and divided them into two subgroups. “Agent” mentioned in the following concepts is Agent Blue and “Opponent” is Agent Red:
i) HowAgentBehaves concepts which are Correct:
– Agent picks high score.
– Agent doesn’t pick high score.
– Agent doesn’t consider opponent’s moves.
ii) HowAgentBehaves concepts which are INcorrect:
– Opponent is random/illogical.
– Agent is random/illogical.
– Agent prolongs game.
– Agent is/can learn
The agent creator considered that these seven concepts reflected on participants’ perceptions of the AI agents’ gameplay, behavior, priorities, strengths, and weaknesses. According to the agent creator, these concepts would cover the required understanding to evaluate a participant’s comprehensive mental model of the AI agents, justifying their integration into the mental model rubric. The agent creator also pointed out that mentioning the correct concepts from subgroup-i indicated participants’ better understanding of the AI agents, while INcorrect concepts from subgroup-ii indicated a lack of understanding. That’s why they assigned positive weights for subgroup-i and negative weights for subgroup-ii in the mental model rubric.
Table 6. Concepts used in the mental model rubric, and the weights/weight ranges Team Game assigned
Concept & Definition Weight Range
Square “X” is miscalculated: Participant says score is wrong/agent miscalculated the score/there are prediction flaws and they specified which particular square they were referring to. [6, 10]
Agent miss-selection: Participant says agent selected the wrong square because it is not the high/highest score (not following its reasoning). 7
Agent chose the wrong action: Participant says agent missed a win opportunity, missed a critical move, or missed an opportunity to block. [11, 15]
Agent picks high score: Participant says agent is picking the square with the highest score or within the top 5. 2
Agent doesn’t pick high score: Participant says agent is not picking the square with the highest score or within the top 5. [6, 10]
Agent doesn’t consider opponent’s moves: Participant says agent does not consider the opponent or consider their effect. [11, 15]
Opponent is random/illogical: Participant says red is dumb, not smart, making weird moves, illogical, random. [-2, -1]
Agent is random/illogical: Participant says agent is generally random, behaving illogically, making weird moves, or doesn’t know what it is doing. [-5, -3]
Agent prolongs game: Participant says agent prolongs/stalls/lengthens the game (e.g., not doing finishing or winning move). [-5, -3]
Agent is/can learn: Participant says agent maybe/is learning. -15
3.4.2. Assignment of Weights: Finding the Weight Ranges
Our team collaborated with the agent creator (Team Game’s lead) to assign weights for the mental model rubric. Team Game’s lead finalized specific weights for certain concepts and a range of weights for others (Table 6). A concept with a range of weights indicates that any values within the range would be appropriate for the mental model rubric, as defined by Team Game’s lead. Our team also provided feedback on possible weights when Team Game’s lead made the decisions. This collaboration ensured our research team and the AI developers both were well-informed about the final weights, aligning with the shared goals of developing the mental model rubric.
From the weight ranges assigned by Team Game’s lead, we got four categories of weight ranges (Table 8). We labeled these categories based on how Team Game’s lead intended the weights to map the importance of the concepts. From these, we derived Table 8’s final weights as follows.
Table 7. Weight categories
Category Weight Range
Low Importance [1, 2], [-2,-1]
Low-Medium Importance [3, 5], [-5, -3]
Medium Importance [6, 10]
High Importance [11, 15]
Table 8. Criteria for weight selection
Criterion Flaws *HAB (Correct) HAB (INcorrect)
Criterion 1: Max within constraints Middle Max Max
Criterion 2: Min within constraints Min Max Min
Criterion 3: INcorrect-Effect-check Max Min Middle
Criterion 4: Flaw-Effect-check Min Max Max
*HAB = HowAgentBehaves
3.4.3. Assignment of Weights: Finalizing the weights
To finalize the weights, we came up with four weighting strategies, described via Table 8’s four criteria. These criteria analyzed the possible effects of the two concept groups within the mental model rubric: i) Flaws, ii) HowAgentBehaves (Correct), and HowAgentBehaves (INcorrect). To check the effects of these concepts, we have considered different combinations of weights (min, middle, max).
For Criteria 1 and 2, we assigned the maximum and minimum weights respectively within their ranges. Criterion 3 varies only in INcorrect weights from Criterion 2 to check if “INcorrect” weight makes much difference. Criterion 4: Flaw-check: varies only Flaws’ weight from Criterion 1 to check if “Flaws” weight makes much difference. As an example of the Min, Middle, and Max weights—the concept “Square X is miscalculated” has a weight range of [6, 10], within this range, we considered its min weight as 6/7, middle weight as 8, and max weight as 9/10.
The rationale behind the weights in the columns of Table 8 were:
For the Flaws concepts, the agent creator emphasized the importance of detecting the flaws as it highlights a participant’s understanding of the agent’s gameplay. However, the agent creator also noted that detecting the flaws was not very difficult, which is why we assigned lower values from their respective weight ranges.
The agent creator reasoned that HowAgentBehaves(correct) provides critical insights into the participant’s overall understanding of the agent’s behavior and recommended assigning higher weights for these concepts. So, we assigned the maximum possible values within their respective weight range in all four criteria.
Lastly, the agent creator wanted the HowAgentBehaves (INcorrect) concepts to account for participants’ incorrect understanding of the AI agents. To maximize the penalty for incorrect understanding of the AI agents, we assigned Max values to these concepts within their ranges.
Table 9 shows the four sets of weights using Table 8’s four criteria/weighting strategies.
Table 9. Exact concept weights used in the four criteria. All are within the ranges specified in Table 6, and were chosen as per the policies in Table 8. C-1 = Criterion 1, C-2 = Criterion 2 etc.
Concept C-1 C-2 C-3 C-4
Square X is miscalculated 8 7 7 7
Agent miss selection 7 7 7 7
Agent chose the wrong action 13 11 11 11
Agent picks high score 2 2 2 2
Agent doesn’t pick high score 10 10 10 10
Agent doesn’t consider opponent’s moves 15 15 15 15
Opponent is random/illogical -2 -1 -1 -2
Agent is random/illogical -5 -3 -4 -5
Agent prolongs game -5 -3 -4 -5
Agent is/can learn -15 -15 -15 -15
4. RQ1 Results: Which participants had better mental model scores, and why?
To answer RQ1, we needed to calculate each participant’s mental score from the concept weights and concept scores, using one of the four criteria from Section 3.1.3’s Table 9—but which one? Fortunately, it turned out that the choice among these four criteria made little difference: the four criteria generated similar participant rankings (Figure 6). For example, 9/10 of the top ten participants by Criterion 1 are also in the top ten participants of Criteria 2, 3, and 4. Likewise, 10/10 of the bottom ten are identical in all four criteria. Indeed, Criteria 2,3, and 4 give almost identical ranking results, as their almost identical overlapping lines show.
Figure 6. Mental model scores (y-axis) for each participant (x-axis) using Criterion 1 (blue) vs. Criterion 2 (red) vs. Criterion 3 (orange) vs. Criterion 4 (purple). Participants are ordered by rank (highest to lowest mental model scores as per Criterion 1). Criteria 2, 3, and 4 rank participants almost identically, and all three rank them similarly to Criterion 1.
We rejected Criterion 1 because its line graph was a little different from the other three criteria, which closely agreed with one another (Figure 6). Of the remaining three, we selected Criterion 4: Flaw-Check for two reasons: first, it avoids overemphasizing “Flaws” concepts and second, because it highlights “HowAgentBehaves” concepts. “Flaws” responses came from the games’ progressing stages, when participants’ mental models were still evolving, so seemed less likely to reflect participants’ final understanding. In contrast, “HowAgentBehaves” responses coming from the end-of-game AAR/AI questions, represented participants’ final mental models by the end of the task, so we assigned maximum weight to the “HowAgentBehaves” concepts. From here on, the term “mental model scores” will refer to the scores calculated using Criterion 4.
4.1. Which participants had better mental model scores?
To compare the two participant groups’ mental model scores, we used a one-tailed t-test. One-tailed t-tests are suitable for detecting effects occurring only in one direction (Frost, 2021), whereas two-tailed tests are for both positive and negative effects. The literature points to one direction only, unanimously reporting GenderMag to improve the user experience of people with diverse problem-solving styles  (Guizani et al., 2022; Vorvoreanu et al., 2019; Burnett et al., 2017; Hilderbrand et al., 2020; Blackmon et al., 2003; Green et al., 2000). Thus, we used Student’s one-tailed t-test to test the following hypotheses, using the mental model score calculations described in Section 3.4.3:
  • Null Hypothesis (H0): The Post-GenderMag participants’ mental model scores are less than or equal to the Original participants’ mental model scores.
  • Alternative Hypothesis (H1): The Post-GenderMag participants’ mental model scores are higher than the Original participants’ mental model scores.
[…]
4.2. Why?—Three Concepts
Why did Post-GenderMag participants develop a better mental model than the Original participants? To find out, we compared each group’s writings about the mental model concepts defined in Section 3.1.3 to identify the key concepts that differentiated the Post-GenderMag participants from the Original participants.
Figure 7 illustrates the differences between the two groups’ written comments about the mental model concepts: concepts above the dashed line were correct statements about how this particular AI worked, and concepts below the dashed line were incorrect statements. (Recall the complete definitions of each concept from Table 6.)
Figure 7. Percentage of Original(n=34) and Post-GenderMag(n=35) participants whose responses included >=>=> =1 statement about that concept. Correct concepts (above the dashed line) increase the mental model score, and incorrect concepts decrease it.
Odds Ratios: The concept is more likely ($>$ 1), equally likely ($=$ 1), or less likely ($<$ 1) to occur in the Post-GenderMag group. *Bold concepts’ Odds Ratios are statistically significant.
Three concepts particularly differentiated the Post-GenderMag participants’ remarks from those of the Original participants: “Agent picks high score”, “Agent doesn’t consider opponent’s moves,” and “Agent is random/illogical”. To measure how much each concept differentiated the two groups, we calculated the odds of Post-GenderMag participants vs. Original participants expressing these concepts; the odds ratios are also in Figure 7. For example, as the figure shows, Post-GenderMag participants were 4.6 times more likely to correctly say that the “Agent doesn’t consider opponent’s moves”; 4 times more likely to correctly say that the “Agent picks high score”; and $\sim$ 4 times less likely to incorrectly say that the “Agent is random/illogical” than the Original participants.
[…]
4.3. The fixes behind the differentiating concepts
These significant differences in participants’ acquisition of the three differentiating concepts raise the question of what particular fixes were tied to these particular differences? We again turned to Post-GenderMag participants’ written comments to see what they had to say about this.
For this purpose, we considered only responses to questions about explanations, because 9 of Team Game’s 13 fixes were changes to explanations. Within those responses, two coders independently coded Post-GenderMag participants’ references to each fix. Their inter-rater reliability (IRR) was 83% with 20% of the data. Given this acceptable level of consensus (Stemler, 2004), the coders then split up the coding of the rest of the data. Figure 8 shows the resulting counts for each fix. As the figure shows, two fixes stood out—Fix-2 and Fix-7.
Figure 8. Total number of times Post-GenderMagparticipants referred to each fix. Fix-2 and Fix-7 stood out—at least 25% of the Post-GenderMag participants referred to these two fixes.
4.3.1. Fix-2: Providing exact win/loss%
In Fix-2 (Figure 3-Bottom), Team Game addressed the issue of incomplete information in the explanations, by providing a detailed breakdown of Agent Blue’s decisions. By highlighting the agent’s predicted win, loss, and draw probabilities, this fix promoted all three differentiating concepts (the agent picking from its own high scores, not using the opponent’s scores, not randomly selecting).
Specifically, whereas in the Original version (Figure 3: Top), participants could only guess at an agent’s exact scoring (win%-loss%) using the graph’s placement of the tiny rectangle of interest, Fix-2 added the option to see the exact win/loss %s with each rectangle of the explanations (Figure  3: Bottom). Upon hovering over a rectangle, the participants could see the exact win%, loss%, and drawif Agent Blue picked the square represented by this rectangle.
Viewing these exact win/loss/draw %s may explain why Post-GenderMag participants were four times more likely to notice “Agent picks high score” (Figure 7). As one Post-GenderMag participant wrote:
P06-PostGM: The AI did appear to attempt to select the squares with the highest probability of winning/highest score ¡Available via Fix-2¿…..
P06-PostGM continued, speculating about how a particular square’s exact win/loss/draw %s affected the probability of winning. Such a specific estimate was made possible by Fix-2’s explicit display of win/loss probabilities:
P06-PostGM (continuing): …..even though I don’t understand how the probability of winning in move 7 with square F1 wasn’t 100%.
Figure 9. Top: (a) Before Fix-2 (Original version): Hovering over a rectangle from the latest data series highlights the rectangles representing the same gameboard square in every data series [^3] Every time agent blue makes a move, the BTW explanation generates a data series containing 36 small rectangles, 1 rectangle for each gameboard square. Where the rectangle lies vertically tells how good the agent thinks moving that square would be. Multiple data series show this info for different decision points.from the previous decision points, but does not show the exact probabilities and does not show which rectangles came from which data series.
Bottom: (b) After Fix-2 (Post-GenderMag version): Hovering over the rectangle now explicitly shows the win%, loss%, and drawof the gameboard square represented by that rectangle.
Showing the agent’s predicted win/lossfrom Fix-2 may have also influenced why Post-GenderMag participants were $4.56$ times more likely to express the concept “Agent doesn’t consider opponent’s moves” compared to the Original participants. This concept is correct—the agent emphasizes its own calculations rather than considering what the opponent agent is doing, as one Post-GenderMag participant pointed out:
P04-PostGM: The AI explanations said the moves had high/highest chances of winning¡Fix-2 showed Win%¿, but it didn’t seem to account for the moves made by red….
Fix-2’s emphasis on the agent’s own calculations also may have helped most Post-GenderMag participants avoid an error that the Original participants made: thinking that the “Agent is random/illogical”, which is not true. Compared to Post-GenderMag participants, Original participants were 4 times more likely to incorrectly assume that Agent Blue was random/illogical. In contrast, Post-GenderMag participants’ responses reflected an understanding of the logic behind Agent Blue’s decisions:
P02-PostGM: BTW allowed me to assess all possibilities the AI was considering & how it percieved it’s win % ¡Available with Fix-2¿.
P68-PostGM: …Chose A2 because of the winand loss¡Available with Fix-2¿. The AI is following the best move for the entire game.
4.3.2. Fix-7: Provided move ranking
Besides Fix-2’s exact win/loss%, Team Game decided to enable participants to compare the top-ranked squares. So with Fix-7, they added an explanation “Top 5 Moves”, which they implemented via Fix-7 (Figure 10). Unlike Fix-2, Fix-7 provided the exact score, which emphasized that Agent Blue usually followed its scores and positions among the top scores. Seeing the exact scores and comparing the top-ranked moves could be the reasons why more Post-GenderMag participants said that “Agent picks high score”:
P07-PostGM: The “scores best-to-worst” and the “top 5 moves” ¡Available with Fix-7¿ help explain it. It shows that it’s ¡Agent Blue¿ picking squares based on their scores.
Participant P39-PostGM also explained how the top 5 moves helped the participant see the agent’s behavior as being purposeful, not random/illogical:
P39-PostGM: Yes - it is clear that X was trying to place moves based on top score (Top5 Moves) ¡Only Post-GenderMag version explicitly showed the top 5 moves¿. It worked in the end.
Figure 10. Fix-7. Post-GenderMag version showed the top 5 moves (red-bordered callout shows at readable size) with their score%, win%, loss%, and draw%.
4.3.3. Causes and the Effects of the fixes
Causes: Both Fix-2 and Fix-7 were about adding more information, driven by insights from Team Game’s GenderMag analyses. As Figure 11 shows Team Game kept pointing to the need for more information because the facet Information Processing Style dominated the GenderMag sessions with Motivations and Attitude Toward Risk following closely behind.
Figure 11. The number of times Team Game mentioned the GenderMag facets during the GenderMag sessions that led to the development of Fix-2 and Fix-7.
[…]
5. RQ2 Results: Prediction Performance and why
5.1. The Prediction Results
To triangulate a different evidence source with the mental model, next, we compared participants’ prediction performance. We calculate participants’ prediction performance using Dodge et al.’s method of calculating “Loss in Value” each per prediction (Koujalgi et al., 2024). Loss in value is a measure of how erroneous a prediction is, we term it PredError in this paper.
PredError’s calculation is:
[…]
where Score is the difference between the AI agent’s predicted $Win\%$ and $Loss\%$. We calculated participants’ PredError for 17 predictions across all three games and their average PredError. A lower PredError indicates better prediction performance.
To compare participants’ prediction performance, we used mixed-design ANOVA for investigating both within-subject (to compare each participant’s 17 predictions) and between-subject (to compare All Original participants’ predictions vs. All Post-GenderMag participants’ predictions) factors. A mixed-design ANOVA can account for both the repeated measures within each participant and the differences between the two groups.
[…]
Figure 12. The PredError for 17 prediction tasks across three games. G=Game, M=Move. For the six prediction tasks not faded—G1M9, G1M11, G2M6, G2M8, G3M7, and G3M13, Post-GenderMagand Originalparticipants’ PredError seemed to differ. Lower PredError indicates higher prediction accuracy.
5.2. Why the Difference? Explanations as both Friend and Foe
But, why did PredError flip-flop between the two groups for those six predictions? Explanations may hold the answer, because almost all differences between the two prototypes were fixes to explanations.
[…]
(2) $\mu($ $\text{Mental Model Score}|\text{Explanation Usage})=\beta_{0}+\beta_{1}\times\text{Explanation Usage}$
where explanation usage is the frequency of each participant’s responses falling under the code “Participant used prior explanation” (Table LABEL:tab:codes_definitions).
Given that the explanations helped participants’ mental models, it seems logical that they would also help participants’ predictions. But what if, in some cases, participants followed the explanations “too closely”? For example, the “Top 5 moves” explanation not only highlights the importance of scores (which helped participants’ mental models), but also shows the moves that were the top 5 for the previous move. If the game remains on a similar trajectory as it was on that previous move, one of those top 5 is likely to be a good choice. However, if Agent Blue’s last move drastically changed the game’s situation, the previous moves will not be good predictions for the next one.
[…]
What this shows us that the Post-GenderMag participants did what Team Game’s new explanations encouraged—basing their predictions on the top 5 moves. This strategy was often effective (the explanation was their “friend”) when the agent’s top moves from the previous move helped in predicting the future move.
For example, in Game 2 Move 6, Post-GenderMag group’s PredError tended to be lower than the Original group (Figure 12). This happened as Post-GenderMag participants selected a move “D3” (Figure 13) less frequently than the Original participants (3/35 Post-GenderMag participants vs 9/34 Original participants). This decision improved Post-GenderMag participants’ prediction accuracy since “D3” was a losing move for Move 6 and the correct prediction was “D4”. Post-GenderMag participants who followed the “top 5 moves” explanation too closely might have avoided “D3” as it was not in the top 5 moves listed (Figure 13). In contrast, Original participants, lacking this explanation, might have considered “D3” a good move based on the gameboard, as it creates multiple options for agent Blue and blocks agent Red.
Figure 13. Gameboard state and the top 5 moves that participants saw before before predicting Game 2: move 6. Here, the explanation “Top 5 Moves” was a friend for Post-GenderMag participants as they avoided selecting “D3” for not being in the top 5 list, which turned out to be a losing move next.
[…]
But, when the top 5 moves did not list advantageous move, the explanation was “foe”. One such example was Game 1 Move 9. For this move, Post-GenderMag participants’ PredError was higher (Figure 12). This might have happened as Post-genderMag participants followed the top 5 moves more than the Original participants. 10/35 Post-GenderMag participants based their predictions on the top 5 moves compared to only 5/34 Original participants. This strategy did not work for the Post-GenderMag group, as “F3”, the correct prediction was outside the top 5. The top 5 moves might have misled the Post-GenderMag participants to overlook the gameboard potential of “F3”(Figure 14) so much that only 8/35 Post-GenderMag participants predicted “F3” compared 20/34 Original participants.
Figure 14. Gameboard state and the top 5 moves that participants saw before predicting Game 1, Move 9. Here the explanation “Top 5 Moves” turned out to be a foe for Post-GenderMag participants. The correct prediction is F3 but they picked mostly from the top 5 moves.
Thus, although the explanations were helpful for Post-GenderMag participants’ mental models, relying on them for predicting the future using the “Top 5 moves” did not always work.
6. RQ3 Results: Inclusivity
RQ3 asks whether the Post-GenderMag prototype provided better inclusivity support. So we investigated if the Post-GenderMag prototype affected participants with “Abi” and “Tim” facets (recall from Table 1) differently than the original version.
Although a facet-by-facet analysis produced raw numbers suggesting improvements for most facet values, no particular facet alone reached significance (Figure 15). However, when considering facet aggregates in individuals by gender, the inclusivity gains became clear.
We begin with two genders—women and men who made up about 93% of our study population, allowing us to perform inferential statistics. For participants from the LGBTQIA* community [^7] LGBTQIA* used based on Scheuerman et al.’s living document (Schmidt and Biessmann, 2020), we then provide a non-statistical representation of their data.
(a) Mental Model Score
(b) PredError
Figure 15. (a): Post-GenderMagAbi (stripes) and Tim (dots) participants had higher mental model scores than their Originalcounterparts. The five facets (from left to right): Info, Learn, Motiv, Risk, and SE (same order in adjacent plots).  (b): Post-GenderMagAbi’s and Tim’s PredError were higher but the difference from the Original group was not significant.
6.1. Inclusivity by Gender
[…]
Figure 16. Post-GenderMagparticipants’ mental model scores were higher than their Originalcounterparts, and the difference was significant for women. The X mark indicates the mean and the horizontal line indicates the median.
[…]
Team Game considered Abi’s facets when making the fixes, so it’s not surprising that the fixes benefitted women as two-thirds ($67\%$) of the women had more Abi facets than Tim (Figure 17). In contrast, men tended more towards Tim ($57\%$) than Abi ($43\%$), yet Team Game’s fixes did not hurt men.
Figure 17. women(n=36) had more Abi facets/problem-solving styles whereas men(n=28) were more spread out but were slightly more Tim than Abi.
Now, what about equity? Equity is “the quality of being fair and impartial” (Press, 2019). Operationally, it would mean men and women achieving the same scores. Figure 18 illustrates the improvement in equity with the Post-GenderMag prototype. In the Original prototype, men and women had a difference of 44 points in their median mental model scores. This gap narrowed to just 24 points in the Post-GenderMag prototype—a 45% improvement in equity. So while improving inclusivity support, Team Game also made the Post-GenderMag prototype more equitable by gender.
Figure 18. Post-GenderMagmen and women show an improvement in equity compared to what Originalmen and women had. The X mark indicates the mean and the horizontal line indicates the median.
Other comparisons such as Post-Gendermag women’s PredError vs Original women’s PredError did not reveal any significant difference. Same thing happened for Post-Gendermag men vs Original men for PredError, mental model scores, and average explanation ratings.
For the comparisons, we used Student’s 2-sided t-tests when all three assumptions of independent samples t-test—1) independent observations, 2) normal distribution, and 3) equal variance were met. The independent observation condition was always true as each group was independent and the participants used the prototype individually with no interaction between them. When the normality distribution failed from a Shapiro-Wilk test, we conducted a non-parametric Wilcoxon Rank-Sum test to compare participants. Lastly, if equal variance failed from an F-test, we conducted a Welch’s t-test for the comparison.
Table 10 reports the GenderMag facet values of the five participants from the LGBTQIA* community. While the dataset is small, we hope it contributes to the growing literature on LGBTQIA* participants (e.g., (Acena and Freeman, 2021; Freeman et al., 2022; Hardy and Vargas, 2019)) and helps future researchers expand our understanding of inclusive design across all gender identities.
Table 10. LGBTQIA* facet values: Each row shows one LGBTQIA* participant’s GenderMag facet values. W = Woman, T = Transgender, NB = Non-Binary, PNS = Prefer Not to State.
PID W NB T PNS SE Motiv Learn Info Risk
P1 Higher Task by Tinkering Selective Tolerant
P11 Higher Tech by Process Selective Averse
P13 Lower Task by Process Comprehensive Tolerant
P46 Lower Task by Process Comprehensive Averse
P59 Higher Tech by Tinkering Comprehensive Tolerant
7. Discussion
7.1. Mental model scores improved but not prediction accuracy, why?
Our results showed that Team Game’s modified explanations helped improve participants’ mental models, but not their prediction accuracy. We wanted to understand why.
7.1.1. Heuristics vs Analytical thinking
The “dual-process theory of cognition” offers a useful lens for understanding the contradictory effects of XAI on our participants. This theory suggests that humans mostly rely on heuristics and mental shortcuts for everyday decisions, and rarely engage in analytical thinking due to its cognitive demand (Buçinca et al., 2021; Kahneman, 2011; Wason and Evans, 1974). Building on this theory, Buçinca et al. developed cognitive forcing functions for XAI to reduce AI overreliance (Buçinca et al., 2021). They posited that explanations primarily help users decide whether and when to follow AI recommendations based on general heuristics, rather than thinking analytically.
This might be true for our participants as well. What if Team Game’s modified explanations only supported participants’ heuristics and could not trigger deeper analytical thinking? We could not confirm this with our data, which could help explain why PostGenderMag participants did not excel in cognitively demanding tasks like predicting the AI agent’s next move.
So, similar to Buçinca et al, would Team Game have benefitted from integrating cognitive forcing functions into the explanations? Buçinca et al.’s study results offer limited optimism. While their cognitive forcing functions reduced AI overreliance, the functions could not improve AI-assisted human decisions (Buçinca et al., 2021).
7.1.2. Declarative vs Procedural knowledge
Post-GenderMag participants achieving significantly higher mental model scores yet no improvement in prediction performance could also stem from what they learn from Team Game’s fixes. Team Game provided declarative knowledge with their fixes, but no procedural knowledge. Declarative knowledge covers facts and information about the “what” of something and procedural knowledge addresses the steps and processes of “how” to do something (Anderson, 2005). In their explanations, Team Game provided “what” the AI considered for making decisions such as win%, loss%, draw%, and score%. However, they did not provide “how” the AI made those decisions (e.g. algorithmic steps, decision paths, or information flows). For Post-GenderMag participants, knowing the “what” led to higher mental model scores. But, they still needed the “how” to improve prediction accuracy.
7.1.3. More ways to fail than to succeed
Also, our participants had more ways to fail. For example, when predicting the first move on an empty gameboard, a participant had only a chance of 1 in 36 to make the correct prediction. We gave them partial credits when they predicted multiple squares and at least one was correct. Still, achieving higher prediction accuracy was difficult. This suggests Team Game’s fixes helped participants excel in basic tasks like understanding the AI agent’s behavior but not so much with advanced tasks like predicting the agent’s next move.
7.2. Abi vs Tim or both?—when making the fixes.
Team Game focused on Abi when finding and fixing the inclusivity bugs and it worked for women, who were predominantly Abi. Readers might wonder if Team Game also needed to consider Tim alongside Abi. We believe the answer is “‘no’ is acceptable but ‘yes’ is better”, and here is why.
Previous GenderMag studies have shown that Abi provides the strongest lens for inclusivity evaluation (Vorvoreanu et al., 2019; Guizani et al., 2022). These studies found more inclusivity bugs with Abi than with Tim, as Abi’s problem-solving styles often required additional inclusivity support. While considering Tim might have usually revealed a few additional inclusivity bugs, the additional time and effort could make the inclusivity evaluation process less feasible. Thus, we recommend following the “Abi-first” approach when the time is too limited to do both.
7.3. Threats to validity
Every study encounters potential challenges to its validity (Wohlin et al., 2000), and ours is no exception. We acknowledge several factors that could limit the generalizability and reliability of our findings.
The first issue is the demographic homogeneity. Although we had 69 participants, most of them were students and came from the same university. This may raise concerns about the broader application of our results. Our participants may not properly reflect the diversity of potential MNK or XAI users.
Moreover, the MNK domain itself presents a potential limitation as this domain is not common in XAI and HAI research. The MNK games Team Game used might not sufficiently replicate real-world scenarios. This could potentially compromise participants’ ability to accurately reason or understand the agents’ behavior. We decided to work with an MNK prototype as people are more familiar with Tic-Tac-Toe which allowed for straightforward judgment on agents’ decisions. Also, MNK gave us the option to ask a wider range of questions related to predictions, flaws, changes, and after-action reviews.
Another concern is that our participants performed proxy tasks. They predicted which square the AI agent would move next, but their predictions did not impact the game. Such proxy tasks may not reflect how users would perform in actual decision-making tasks. As Buçinca et al. showed evaluations of XAI systems with proxy tasks did not predict the results of evaluations with actual decision-making tasks. While our proxy tasks reduced cognitive load and allowed more focus on explanations, they might have limited us from creating immersive and high-stakes scenarios. An important future research would be evaluating if GenderMag-driven fixes can improve users’ performance in actual decision-making tasks.
We did not gather click log data for explanation interactions. Instead, we relied on participants’ mentions of using the explanations as an engagement indicator. Maintaining a click log could have offered more dependable insights on explanation usage. While we have screen recordings from each session, determining the time spent specifically on explanations would be difficult and time-consuming.
Addressing such threats would require additional studies using diverse empirical methods. Those methods have to make sure the findings apply widely to various XAI systems and participant demographics.
8. Conclusion
This paper investigated if inclusivity fixes can improve users’ mental models of the AI system. Our setting was an XAI prototype where participants observed two AI agents sequentially play a modified version of Tic-Tac-Toe. We report the results from our between-subject study where the “Original” group used the original version of the prototype and the “Post-GenderMag” group used the version with the inclusivity fixes:
  • RQ1: Participants who used the Post-GenderMag version of the prototype demonstrated significantly better mental models than the Original version participants (Section 4).
  • RQ2: There was no significant difference in the two groups of participants’ prediction performance (Section 5).
  • RQ3: Inclusivity fixes made the Post-GenderMag prototype more inclusive for women without hurting men and improved equity by reducing the gender gap (Section 6).
These results indicate the impacts of inclusivity fixes in improving users’ mental models and effectively increasing equity and inclusion within XAI systems. Our study highlights the potential of problem-solving style-centric inclusive design approaches to create more accessible and user-centric XAI prototypes, benefiting diverse users’ understanding of AI.
P37-PostGM: The BTW ¡Best to Worst explanation¿ and top 5 moves score ¡available in the Post-GenderMag version only¿ helped me gain a better understanding of why the AI did the things it did.
Appendix
Table 11. Codes and definitions
Code Definition
Prediction Why
Participant used current game state Participant drew on aspects of the current game state, such as gameboard, move placement, sequence orientation, win/loss position/state (condition of gameboard).
Participant used prior actions Participants drew on something they’d seen the agent do (games/moves).
Participant used prior explanation Participants drew on something they’d seen in the previous explanation(s) including the explanations from a previous game.
Participant predicted next explanation Participants hypothesized what the explanation would look like after the agent took the action they predicted.
Participant said agent was random or baseless Participant stated something about the agent that seemed random/wasn’t made for a “good” reason.
Participant stated what they’d do Participants placed themselves as a pilot, directing what they want to attack.
Participant is randomly guessing Participant gives a random guess, particularly for an empty gameboard or early game.
Flaws
Yes failure Participant detected an issue, manifested an incorrect output i.e., wrong move, value too high/low, missing opportunity, missing an opportunity to prevent a loss.
Yes fault Participant identified why the agent performed poorly or something wrong about how the agent processed its input, or agent is focusing too much on its own moves.
General miscalculation Participant says or perceives the score is wrong/agent miscalculated the score/there are prediction flaws but they did not specify which particular square they were referring to.
Square “X” is miscalculated Participant says the score is wrong/agent miscalculated the score/there are prediction flaws and they specified which particular square they were referring to.
Agent miss-selection Participant says the agent selected the wrong square because it is not the high/highest score (not following its reasoning).
Agent chose the wrong action Participant says there is a lost win opportunity, missed a critical move, or missed the opportunity to block.
General flaw Participant only mentioned that there is a flaw but did not specify why it’s a flaw.
Strategy Should be Participant says some general idea about how the agent SHOULD play the game or suggestions that the agent should follow.
Agent should pick high score Participant says the agent should use the highest / high score or top 5.
Agent should do location-based Strategy Participant says the agent should pick based on the gameboard, past moves, orientation, or should continue the sequence to 4 in a row, prioritize established sequence unless under threat, take winning move, block opponent’s move, follow offensive or defensive strategy.
Agent should create multiple options Participant says the agent should create multiple options, e.g., 2 2 X in a row.
Agent should follow other strategy Participant says the agent should follow another strategy such as changing how it calculates, reduce the number of moves, and/or consider more options/data (e.g., consider opponent’s moves).
HowAgentBehaves Participant’s perception of the agent’s behavior, such as how the agent plays the game, how it behaves during the game, what its priority, strengths and weaknesses are.
Agent picks high score Participant says the agent is picking the square with the highest score or within the top 5.
Agent doesn’t pick high score Participant says the agent is not picking the square with the highest score or within the top 5.
Agent does location-based Strategy Participant says the agent picks based on the gameboard, past moves, orientation, or will continue the sequence to 4 in a row, prioritize established sequence unless under threat, take winning move, block opponent’s move, follow offensive or defensive strategy.
Agent does not do location-based Strategy Participant says the agent does not pick based on the gameboard, past moves, orientation, continue the sequence to 4 in a row, prioritize established sequence unless under threat, take winning move, block opponent’s move, follow offensive or defensive strategy.
Agent doesn’t consider opponent’s moves Participant says the agent does not consider the opponent or consider their effect.
Agent does make multiple options Participant says the agent makes multiple options, e.g., 2 2 X in a row.
Agent thinks ahead Participant says explicitly that the agent can think ahead.
Agent doesn’t think ahead Participant says that the agent does not think ahead.
Opponent is random/illogical Participant says red is dumb, not smart, making weird moves, illogical, random.
Agent is random/illogical Participant says generally the agent is random, behaving illogically, making weird moves, or doesn’t know what it is doing.
Agent prolongs game Participant says the agent prolongs/stalls/lengthens the game (e.g., not doing finishing or winning move).
Agent is/can learn Participant says the agent maybe/is learning.
Participant has uncertainty/confusion Participant said they are unsure/confused about something with respect to how the agent works (possibly speculatively).
Participant has a question Participant explicitly asks a question about the agent, game, or explanations and may want more information.
Starting first has advantages Participant says that going first has advantages or going second has fewer advantages towards winning the game.
Start doesn’t matter Participant says that starting moves do not matter.
Score should be symmetrical Participant says scores should be symmetrical with respect to the gameboard.
Table 12. All of Team J’s Fixes, why they fixed them (problem-solving styles impacted), and the changes made for the fixes
Fix-ID & Fix Name Fixed-Why + Problem Solving Styles(PSS) Changes Made for the Fixes
Fix-1: Blurb of Instructions ”Attitude toward technology is by process,…look for documentation..” PSS: Info, Risk. – Added instructions on getting started. – Added summary of each explanation. – Added minimize button for instructions.
Fix-2: Exact values and Exact actions ”She is not clear on what she should do…..will not take this step.” PSS: Risk, Learn, Info, SE, Motiv. – Showed exact win%, loss%, and drawfor every rectangle in the BTW, STT, and OTB explanations using tooltips. – Added tooltips that informed about the exact actions offered by the next and the prior move buttons.
Fix-3: Buttons Design and Colors ”The game info…gives the hint….needs further indications.” PSS: Motiv, Info. – Made buttons look more like buttons. – Increased color contrast between the button and the instructions background. – Used undo icon with the prior move button.
Fix-4: Info Button “Not sure if Abi knows….what type of info it would provide.” PSS: SE, Learn. – Added circular info button with each explanation that gives a short summary of the explanation.
Fix-5: Colors “Hovering does….This falls into her learning style..” PSS: SE, Learn. – Changed the color of the interactive elements. – Increased the contrast between the interactive element and the interface background.
Fix-6: Labels and Titles “She sees new information…needs to understand what….trying to provide.” PSS: Motiv, Info. – Changed labels for STT x-axis: Move Number instead of Time. – Changed the BTW and STT y-axis: 100% win and 100% loss instead of just win and loss. – Added a title bar “Spot the Broken AI”.
Fix-7: Top 5 Moves “…but they may not be sure how to reason with this information.” PSS: Risk, Learn, Info, Motiv. – Added “Top 5” moves with their score, win, loss, and draw percentages.
Fix-8: Tying Arrows “..no update for the other player…she may consider….something wrong.” PSS: Info, Risk, SE. – Added arrows on BTW, STT, OTB that point toward the rectangles representing the scores of the most recent move. – These arrows appeared for a few seconds when the user clicked the next/prior move button.
Fix-9: Game Log “…nothing has changed….so Abi might be confused.” PSS: Risk. – Added a game log beside the move rank list. – Added info on all the moves in the log. – Added a scroll bar and a separator between the game log and the top 5 moves.
Fix-10: BTW Legend “The relationship between…board and the BTW…isn’t clear.” PSS: SE, Risk. – Added legend to the top right for btw. – Changed the coloring of the move lines. Darker color for more recent moves. – Made the legend interactive.
Fix-11: Text and Prompts “Nothing is clearly stating that the game is over…” PSS: SE. – Added a text message at the beginning shows the game is starting. – Added a text message at the end that shows the game has ended. – Alongside the last move’s number added who won. – Removed who moves next for the last move. – Mentioned the game number with the move, e.g., Game 1 Move 1.
Fix-12: Chart Highlight changes “When highlighting…, there are multiple….This causes confusion for Abi..likes to have a complete understanding…” PSS: Info, SE, Learn. – Added the same highlighting on OTB for the winning sequence. – Added a thick vertical bar on STT indicates that this is the last move. – Changed BTW interaction to just highlight the rectangle from the current data series.
Fix-13: Game History “Because of IP/Risk, …rethink what they’re doing before moving on.” PSS: Info, SE, Risk. – Provided a drop-down game history at the top. – Provided a summary of what happened in the earlier games inside the game history drop-down.
Fix-14: Move History “Abi will want to spend more time, because with the prior move…what it’s telling her.” PSS: Info, SE, Risk. – Team J wanted to have the feature to go back to any move from an older game. We have not implemented it.
Fix-15: Zoom in/out “Not sure that Abi would get the information…given her information processing style. PSS: Info, SE. – Team J wanted to have a zoom-in/out feature for the explanation charts. We have not implemented it.
9 of the 15 fixes: Fix-1, 2, 4, 5, 6, 7, 8, 10, and 15 were targeted to improve the explanations.